OpenStudy (anonymous):

Twelve and seventeen-eighteenths minus 4 and two-thirds is equal to what number? 8 and five-sixths 8 and five-thirteenths 8 and five-eighteenths 3 and one-sixteenth

OpenStudy (diyadiya):

\[12 \frac{17}{18}- 4 \frac{2}{3}\]

OpenStudy (diyadiya):

First convert it to proper fraction

OpenStudy (diyadiya):

\[12 \frac{17}{18}= \frac{12 \times 18+17}{18}=233/18\]

OpenStudy (diyadiya):

Can you convert 4 2/3 ?

OpenStudy (anonymous):

is it 

OpenStudy (diyadiya):

No

OpenStudy (anonymous):

owe

OpenStudy (diyadiya):

\[4 \frac{2}{3}= \frac{4 \times 3+2}{3}= \frac{14}{3}\] Now do \[\frac{233}{18}- \frac{14}{3}\]

OpenStudy (anonymous):

233-14=219 18-8=15

OpenStudy (diyadiya):

No ,you can subtract fractions with different denominator ,you should make it to common denominator\[\frac{233}{18}-\frac{14 \times 6}{3 \times 6}= \frac{233}{18}-\frac{84}{18}= \frac{149}{18}= 8 \frac{5}{18}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks